A vessel contains oil of density 0.8 cm-3 over mercury of density 13.6 g cm-3. A homogeneous sphere floats with ¼ th of its volume immersed in mercury and the rest ¾ th in oil. The density of the material of the sphere is ?

Grade:12th pass

1 Answers

Arun
4 years ago
 

The total weight is balanced by the displaced volumess of the two different fluids. Let it,s Vol is V and the density is n. Then the equilibrium condition should be

 Vng=V/2(0.8)g+V/2(13.6)g

or n=0.4+6.8=7.2 gm/cc(ans)